OpenStudy (anonymous):

Hello, can someone help us. My son has a math question that we just cannot figure out: 1/2XN=15

OpenStudy (anonymous):

Is it (1/2) * (xn) = 15 or 1/(2xn) = 15

OpenStudy (anonymous):

And does it want you to solve for x?

OpenStudy (anonymous):

I am sorry for the confusion it should read 1/2 * N=15 I think they want to solve for .